please solve and show work, thanks
6x^2-11x=2

Answer by bubblz190(11)

6x^2 - 11x = 2
6x^2 - 11x - 2 = 0
Now you would put that equation into the quadratic formula.
x = (-b +- sqrt( b^2-4*a*c ))/(2*a)
a=6 b=-11 c=-2
x = (11 +- sqrt(121 + 48))/12
x = (11 +- 13)/ 12
x = (2, -1/6)